# Test-time Adaptive Hierarchical Co-enhanced Denoising Network for Reliable Multimodal Classification

### v1

@@ III-C1 Prior Estimation
III-C2 Sample-Adaptive Noise Expert
Modality-specific and cross-modality noise experts are then introduced to perform noise removal on each sample under the guidance of the estimated priors. Specifically, we introduce $M$ modality-specific noise experts $E_{s}^{1},\ldots,E_{s}^{M}$} and $M$ cross-modality noise experts $E_{c}^{1},\ldots,E_{c}^{M}$}, each corresponding to one modality. For a multimodal sample $x_{i}=\{x_{i}^{m}\}_{m=1}^{M}$ with globally denoised representation $h_{i}=\{h_{i}^{m}\}_{m=1}^{M}$, each modality representation $h_{i}^{m}$ is fed into $E_{s}^{m}$ and $E_{c}^{m}$ to learn masks $w_{s,i}^{m}$ and $w_{c,i}^{m}$ that capture feature-wise informativeness relevant to modality-specific and cross-modality noise, respectively. These masks are then used to remove the modality-specific and cross-modality noise components $n_{s,i}^{m},n_{c,i}^{m}$ from $h_{i}^{m}$ and yield the corresponding denoised features $\hat{h}_{s,i}^{m},\hat{h}_{c,i}^{m}$. The calculation in $E_{s}^{m}$ can be formulated as:
@@ III-C2 Sample-Adaptive Noise Expert
III-C3 Confidence-Aware Asymmetric Slack Alignment
@@
$\mathcal{L}_{\mathrm{saca}}=\mathcal{L}_{\mathrm{nll}}^{s}+\mathcal{L}_{\mathrm{nll}}^{c}$ (12)
$=-\sum_{m=1}^{M}\log p^{m}_{h}(\hat{h}_{s}^{m})-\sum_{m=1}^{M}\sum_{m^{\prime}>m}^{M}\log p^{m-m^{\prime}}_{h}(\hat{h}_{c}^{m}-\hat{h}_{c}^{m^{\prime}}).$
@@
It is worth noting that, in $\mathcal{L}_{\mathrm{nll}}^{c}$, unlike existing methods that remove cross-modality noise by maximizing inter-modality feature similarity, we instead constrain the feature discrepancies across modalities within a reasonable range. This range accounts for the inherent informative information differences across modalities, which is learned by the cross-modality discrepancy distribution $p_{h}^{m-m^{\prime}}$. We refer to this strategy as slack alignment, which enables the removal of cross-modality noise while preserving more modality-specific complementary information, thereby facilitating more reliable multimodal representation learning. Moreover, considering confidence differences across modalities under modality-specific noise, we implement the slack alignment in an asymmetric manner directed by modality confidence. This encourages low-confidence modality features to align toward high-confidence ones, imposing stronger rectification on them, instead of rectifying uniformly across all modalities. Specifically, the update of the parameters in cross-modality noise expert $E_{c}^{m}$ under the guidance of modality confidence can be expressed as:
@@
$\theta_{c}^{m}\leftarrow\theta_{c}^{m}-\eta\cdot\frac{1}{|B|}\sum_{i}^{|B|}\neg\mathrm{conf}^{m}_{i}\cdot\nabla_{\theta_{c}^{m}}\mathcal{L}_{\mathrm{nll}}^{c}(\hat{h}_{c}^{m};\theta_{c}^{m}).$ (13)
@@
$\theta_{c}^{m}$ denotes the network parameters of the cross-modality noise expert $E_{c}^{m}$, $\eta$ is the learning rate, $B$ is a random mini-batch, and $\nabla_{\theta_{c}^{m}}\mathcal{L}_{\mathrm{nll}}^{c}$ represents the gradient of $\theta_{c}^{m}$ with respect to $\mathcal{L}_{\mathrm{nll}}^{c}$. $\neg\mathrm{conf}^{m}_{i}$ is a factor negatively correlated with the confidence of modality $m$ feature $h^{m}_{i}$ and can be dynamically computed during training as follows:
$\neg\mathrm{conf}_{i}^{m}=\frac{\exp(1-\mathrm{tanh}(\mathrm{c}_{i}^{m}))}{\sum_{m^{\prime}=1}^{M}\exp(1-\mathrm{tanh}(\mathrm{c}_{i}^{m^{\prime}}))},$ (14)
@@
where $\mathrm{c}_{i}^{m}=p^{m}(h_{i}^{m})$ is the confidence of sample feature $h_{i}^{m}$ of modality $m$, estimated as its likelihood under under the modality-$m$ feature prior $p^{m}$. In this way, we amplify the update gradients of cross-modality noise experts responsible for low-confidence modality representations while reducing those for high-confidence ones. This encourages the model to focus on mining and applying stronger alignment rectification to low-confidence modalities with pronounced modality-specific noise, while applying weaker rectification to high-confidence ones with less modality-specific noise. Consequently, cross-modality noise is more effectively mitigated in the presence of modality-specific noise, promoting the joint removal of both noise types.
III-D Test-Time Cooperative Enhancement
@@
ASSA and SACA establish a framework for jointly removing modality-specific and cross-modality noise at both global and individual levels to achieve robust learning. Building upon this framework, we introduce Test-Time Cooperative Enhancement (TTCE), which performs adaptive model updates driven by sample noise in a label-free manner to enhance the model’s generalization to unseen noise. Specifically, TTCE is designed as a cooperative enhancement mechanism between ASSA and SACA conditioned on unseen noise.
@@
Considering that instance-level representations are more sensitive to noise variations than global distributions, TTCE first enhance global-level denoising of ASSA using the learned instance-level noise from SACA. Specifically, we incorporate instance-level modality-specific and cross-modality noise components $n_{s}^{m}=\{n_{s,i}^{m}\}_{i=1}^{N}$ and $n_{c}^{m}=\{n_{c,i}^{m}\}_{i=1}^{N}$ from SACA into the globally denoised modality features $h^{m}$ from ASSA and define a reconstruction loss:
@@
$\Psi^{m}(\cdot)$ denotes the decoder corresponding to modality $m$, and $\|\cdot\|_{\mathrm{F}}^{2}$ denotes the Frobenius norm. Minimizing $\mathcal{L}_{\mathrm{re}}$ enables ASSA to explicitly account for more accurate and finer-grained unseen-noise information captured at the instance level, yielding more robust globally denoised modality representations $h^{m\uparrow}$ to unseen noise. The process of enhancing $h^{m}$ to $h^{m\uparrow}$ via $\mathcal{L}_{\mathrm{re}}$ can be expressed as:
$h^{m\uparrow}$ $=h^{m\uparrow}-\eta\frac{\partial\mathcal{L}_{\mathrm{re}}}{\partial h^{m\uparrow}}$
$=h^{m\uparrow}-2\eta J_{\Psi^{m}}(I_{t}^{m})^{\top}(\Psi^{m}(I_{t}^{m})-x^{m}),$ (16)
where $h^{m\uparrow}$ is initialized as $h^{m}$ at the beginning of the update process. $I_{t}^{m}=\tilde{z}_{t}^{m}+n_{s}^{m}+n_{c}^{m}$. $J_{\Psi^{m}}$ represents the Jacobian matrix of the decoder $\Psi^{m}$ evaluated at the input. $\eta$ is the learning rate.
@@
$\mu^{m\uparrow}$ $=\alpha\cdot\mu^{m}+(1-\alpha)\cdot\Delta\mu^{m\uparrow},$
$\Sigma^{m\uparrow}$ $=\beta\cdot\Sigma^{m}+(1-\beta)\cdot\Delta\Sigma^{m\uparrow}.$ (17)
$\mu^{m\uparrow}$ and $\Sigma^{m\uparrow}$ are the mean and covariance of the updated modality-wise distribution $p^{m}$, denoted as $p^{m\uparrow}$. $\alpha$ and $\beta$ are two hyperparameters. Accordingly, the cross-modality discrepancy distribution $p^{m-m^{\prime}}(m,m^{\prime}\in[1,M],m\neq m^{\prime})$ is updated to $p^{m-m^{\prime}\uparrow}\sim\mathcal{N}(\mu^{m-m\prime\uparrow},\Sigma^{m-m^{\prime}\uparrow})$, with $\mu^{m-m\prime\uparrow}$ and $\Sigma^{m-m^{\prime}\uparrow}$ derived from $\mu^{m\uparrow}$, $\mu^{m\prime\uparrow}$, $\Sigma^{m\uparrow}$, and $\Sigma^{m\prime\uparrow}$ through Eq. (9). The enhanced priors $p^{m\uparrow}$, $p^{m\prime\uparrow}$, and $p^{m-m\prime\uparrow}$ are then applied in $\mathcal{L}_{\mathrm{saca}}$ defined in Eq. (12) to improve the instance-level noise removal of the noise experts against unseen noise, which can be formulated as:
@@
$\hat{h}_{c}^{m\uparrow}=\hat{h}_{c}^{m\uparrow}-\eta\frac{\partial\mathcal{L}_{\mathrm{saca}}}{\partial\hat{h}_{c}^{m\uparrow}}$ (18)
$=\hat{h}_{c}^{m\uparrow}-\eta\sum\nolimits_{m^{\prime}\neq m}^{M}{(\Sigma^{m-m^{\prime\uparrow}})^{-1}(\hat{h}_{c}^{m\uparrow}-\hat{h}_{c}^{m\prime\uparrow}-\mu^{m-m^{\prime\uparrow}})}.$
Therefore, through multiple iterations of the above ASSA–SACA cooperative enhancement, TTCE progressively updates the model to adapt to unseen noise. For clarity, the overall inference procedure of TAHCD with TTCE is summarized in Algorithm 1, where $E$ denotes the number of cooperative enhancement iteration between ASSA and SACA and is analyzed in Sec. IV-G.
@@
Input: Multimodal input $X=\{\{x_{i}^{m}\}_{i=1}^{N}\}_{m=1}^{M}$, trained ASSA and SACA.
@@
10: Fuse $\{\hat{h}_{s}^{m}\}_{m=1}^{M}$, $\{\hat{h}_{c}^{m}\}_{m=1}^{M}$ using Eq. (19) for prediction.
III-E Training and Inference
Finally, we fuse the multimodal features produced by sample-adaptive noise experts after TTCE, weighted by their respective confidence scores. Specifically, the confidence score is first estimated as the likelihood of the input feature vector under the corresponding modality-wise distribution: $\mathrm{conf}_{s}^{m\uparrow}=p^{m\uparrow}(\hat{h}_{s}^{m\uparrow}),\mathrm{conf}_{c}^{m\uparrow}=p^{m\uparrow}(\hat{h}_{c}^{m\uparrow})$. We then normalize the output confidences of all experts to obtain weights $\{\overline{\mathrm{conf}}_{s}^{m\uparrow}\}_{m=1}^{M},\{\overline{\mathrm{conf}}_{c}^{m\uparrow}\}_{m=1}^{M}$, which are used to fuse $\{h_{s}^{m\uparrow}\}_{m=1}^{M},\{h_{c}^{m\uparrow}\}_{m=1}^{M}$:
@@ III-E Training and Inference
IV Experiments
In this section, we conduct extensive experiments to validate the robustness and generalization ability of the proposed Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD) under multimodal noise. Experimental Setups are first introduced in Sec. IV-A. Comparisons with state-of-the-art methods and ablation studies are then presented in Secs. IV-B and IV-C, respectively. Discussions on different model components, together with parameter analyses and additional investigations, are detailed in Secs. IV-DIV-G.

### v2

@@ 3.3.1 Prior Estimation
3.3.2 Sample-Adaptive Noise Expert
Modality-specific noise experts and cross-modality noise experts are then introduced to perform noise removal on each sample under the guidance of the estimated priors. Specifically, we introduce $M$ modality-specific noise experts $E_{s}^{1},\ldots,E_{s}^{M}$} and $M$ cross-modality noise experts $E_{c}^{1},\ldots,E_{c}^{M}$}, each corresponding to one modality. For a multimodal sample $x_{i}=\{x_{i}^{m}\}_{m=1}^{M}$ with globally denoised representation $h_{i}=\{h_{i}^{m}\}_{m=1}^{M}$, each modality representation $h_{i}^{m}$ is fed into $E_{s}^{m}$ and $E_{c}^{m}$ to learn masks $w_{s,i}^{m}$ and $w_{c,i}^{m}$ that capture feature-wise informativeness relevant to modality-specific and cross-modality noise, respectively. These masks are then used to remove the modality-specific and cross-modality noise components $n_{s,i}^{m},n_{c,i}^{m}$ from $h_{i}^{m}$ and yield the corresponding denoised features $\hat{h}_{s,i}^{m},\hat{h}_{c,i}^{m}$. The calculation in $E_{s}^{m}$ can be formulated as:
@@ 3.3.2 Sample-Adaptive Noise Expert
3.3.3 Confidence-Aware Asymmetric Slack Alignment
@@
$\mathcal{L}_{\mathrm{saca}}=\mathcal{L}_{\mathrm{nll}}^{s}+\mathcal{L}_{\mathrm{nll}}^{c}=\left[-\sum_{m=1}^{M}\log p^{m}_{h}(\hat{h}_{s}^{m})\right]+\left[-\sum_{m=1}^{M}\sum_{m^{\prime}>m}^{M}\log p^{m-m^{\prime}}_{h}(\hat{h}_{c}^{m}-\hat{h}_{c}^{m^{\prime}})\right].$ (12)
@@
It is worth noting that, in $\mathcal{L}_{\mathrm{nll}}^{c}$, unlike existing methods that remove cross-modality noise by maximizing similarity across modalities, we instead constrain the feature discrepancies $(\hat{h}_{c}^{m}-\hat{h}_{c}^{m^{\prime}})$ across modalities within a reasonable range. This range accounts for the inherent informative information differences across modalities, which is learned by the cross-modality discrepancy distribution $p_{h}^{m-m^{\prime}}$. We refer to this strategy as slack alignment, which enables the removal of cross-modality noise while preserving more modality-specific complementary information, thereby facilitating more reliable multimodal representation learning. Moreover, considering confidence differences across modalities under modality-specific noise, we implement the slack alignment in an asymmetric manner directed by modality confidence. This encourages low-confidence modality features with more modality-specific noise to align toward high-confidence ones, imposing stronger rectification on them, instead of rectifying uniformly across all modalities. Specifically, the update of the parameters in cross-modality noise expert $E_{c}^{m}$ under the guidance of modality confidence can be expressed as:
@@
$\theta_{c}^{m}\leftarrow\theta_{c}^{m}-\eta\cdot\frac{1}{|B|}\sum_{i}^{|B|}\neg conf^{m}_{i}\cdot\nabla_{\theta_{c}^{m}}\mathcal{L}_{\mathrm{nll}}^{c}(\hat{h}_{c}^{m};\theta_{c}^{m}).$ (13)
@@
$\theta_{c}^{m}$ denotes the parameters of the cross-modality noise expert $E_{c}^{m}$ corresponding to modality $m$. $\eta$ is the learning rate, $B$ is a random mini-batch, and $\nabla_{\theta_{c}^{m}}\mathcal{L}_{\mathrm{nll}}^{c}$ is the gradient of $\theta_{c}^{m}$ with respect to $\mathcal{L}_{\mathrm{nll}}^{c}$. $\neg conf^{m}_{i}$ is a factor negatively correlated with the confidence $c_{i}^{m}$ of modality-$m$ feature $h^{m}_{i}$ and is dynamically computed during training as follows:
$\neg conf_{i}^{m}=\frac{\exp(1-\mathrm{tanh}(c_{i}^{m}))}{\sum_{m^{\prime}=1}^{M}\exp(1-\mathrm{tanh}(c_{i}^{m^{\prime}}))},$ (14)
@@
where $c_{i}^{m}=p^{m}(h_{i}^{m})$ is estimated as the likelihood of $h_{i}^{m}$ under under the modality-$m$ feature prior $p^{m}$. In this way, the update gradients of cross-modality noise experts are amplified for low-confidence modality and reduced for high-confidence ones. This encourages the model to focus on mining and applying stronger alignment rectification to lower-confidence modalities with more pronounced modality-specific noise. Consequently, the model achieves more reliable cross-modality noise mitigation under modality-specific noise, enhancing the joint removal of both noise types.
3.4 Test-Time Cooperative Enhancement
@@
ASSA and SACA establish a framework for jointly removing modality-specific and cross-modality noise at both global and individual levels to achieve robust learning. Building on this framework, we introduce Test-Time Cooperative Enhancement (TTCE), which performs label-free adaptive updates guided by sample noise to improve generalization to unseen noise through cooperative enhancement of ASSA and SACA.
@@
Considering that instance-level representations are more sensitive to noise variations than global distributions, noise information learned at the instance level can capture richer cues about unseen noise than that at the global level. Therefore, TTCE proposes to first leverage instance-level noise learned by SACA to enhance global-level denoising of ASSA under unseen noise. Specifically, we incorporate instance-level modality-specific and cross-modality noise components $n_{s}^{m}=\{n_{s,i}^{m}\}_{i=1}^{N}$ and $n_{c}^{m}=\{n_{c,i}^{m}\}_{i=1}^{N}$ from SACA into the globally denoised modality features $h^{m}$ from ASSA and define a reconstruction loss:
@@
$\Psi^{m}(\cdot)$ denotes the decoder corresponding to modality $m$, and $\|\cdot\|_{\mathrm{F}}^{2}$ denotes the Frobenius norm. Minimizing $\mathcal{L}_{\mathrm{re}}$ enables ASSA to explicitly account for more accurate and finer-grained unseen-noise information captured at the instance level, yielding more reliable globally denoised modality representations $h^{m\uparrow}$ to unseen noise. The process of enhancing $h^{m}$ to $h^{m\uparrow}$ via minimizing $\mathcal{L}_{\mathrm{re}}$ can be equivalently expressed as:
$h^{m\uparrow}$ $=h^{m\uparrow}-\eta\frac{\partial\mathcal{L}_{\mathrm{re}}}{\partial h^{m\uparrow}}=h^{m\uparrow}-2\eta J_{\Psi^{m}}(I^{m})^{\top}(\Psi^{m}(I^{m})-x^{m}),$ (16)
where $h^{m\uparrow}$ is initialized as $h^{m}$ at the beginning of the update process. $I^{m}=h^{m}+n_{s}^{m}+n_{c}^{m}$. $J_{\Psi^{m}}$ represents the Jacobian matrix of the pre-trained decoder $\Psi^{m}$ evaluated at the input. $\eta$ is the updating rate.
@@
$\mu^{m\uparrow}$ $=(1-\alpha)\cdot\mu^{m}+\alpha\cdot\Delta\mu^{m\uparrow},$
$\Sigma^{m\uparrow}$ $=(1-\alpha)\cdot\Sigma^{m}+\alpha\cdot\Delta\Sigma^{m\uparrow}.$ (17)
$\mu^{m\uparrow}$ and $\Sigma^{m\uparrow}$ are the mean and covariance of the updated modality-wise distribution $p^{m}$, denoted as $p^{m\uparrow}$. $\alpha\in[0,1]$ is a hyperparameter, which will be analyzed in Sec. 4.8. Accordingly, the cross-modality discrepancy distribution $p^{m-m^{\prime}}(m,m^{\prime}\in[1,M],m\neq m^{\prime})$ is updated to $p^{m-m^{\prime}\uparrow}\sim\mathcal{N}(\mu^{m-m\prime\uparrow},\Sigma^{m-m^{\prime}\uparrow})$, with $\mu^{m-m\prime\uparrow}$ and $\Sigma^{m-m^{\prime}\uparrow}$ derived from $\mu^{m\uparrow}$, $\mu^{m\prime\uparrow}$, $\Sigma^{m\uparrow}$, and $\Sigma^{m\prime\uparrow}$ through Eq. (9). The enhanced priors $p^{m\uparrow}$, $p^{m\prime\uparrow}$, and $p^{m-m\prime\uparrow}$ are then applied in $\mathcal{L}_{\mathrm{saca}}$ defined in Eq. (12) to improve the instance-level noise removal of the noise experts against unseen noise, which can be equivalently expressed as:
@@
$\hat{h}_{c}^{m\uparrow}=\hat{h}_{c}^{m\uparrow}-\eta\frac{\partial\mathcal{L}_{\mathrm{saca}}}{\partial\hat{h}_{c}^{m\uparrow}}=\hat{h}_{c}^{m\uparrow}-\eta\sum\nolimits_{m^{\prime}\neq m}^{M}{(\Sigma^{m-m^{\prime\uparrow}})^{-1}(\hat{h}_{c}^{m\uparrow}-\hat{h}_{c}^{m\prime\uparrow}-\mu^{m-m^{\prime\uparrow}})}.$ (18)
Therefore, through multiple iterations of the above cooperative enhancement between ASSA and SACA, TTCE progressively updates the model to adapt to unseen noise. For clarity, the overall inference procedure of TAHCD with TTCE is summarized in Algorithm 1, where $E$ denotes the number of cooperative enhancement iteration between ASSA and SACA and will be analyzed in Sec. 4.6.3.
@@
Input: Multimodal input $X=\{\{x_{i}^{m}\}_{i=1}^{N}\}_{m=1}^{M}$, trained TAHCD.
@@
10: Fuse $\{\hat{h}_{s}^{m\uparrow}\}_{m=1}^{M}$, $\{\hat{h}_{c}^{m\uparrow}\}_{m=1}^{M}$ using Eq. (19) for prediction.
3.5 Training and Inference
Finally, we fuse the multimodal features $\{h_{s}^{m\uparrow}\}_{m=1}^{M},\{h_{c}^{m\uparrow}\}_{m=1}^{M}$ produced by sample-adaptive noise experts after TTCE, weighted by their respective confidence scores. Specifically, the confidence score is first estimated as the likelihood of the feature vector under the corresponding modality-wise distribution: $\mathrm{conf}_{s}^{m\uparrow}=p^{m\uparrow}(\hat{h}_{s}^{m\uparrow}),\mathrm{conf}_{c}^{m\uparrow}=p^{m\uparrow}(\hat{h}_{c}^{m\uparrow})$. We then normalize the confidence scores of all experts to obtain weights $\{\overline{\mathrm{conf}}_{s}^{m\uparrow}\}_{m=1}^{M},\{\overline{\mathrm{conf}}_{c}^{m\uparrow}\}_{m=1}^{M}$, which are used to fuse $\{h_{s}^{m\uparrow}\}_{m=1}^{M},\{h_{c}^{m\uparrow}\}_{m=1}^{M}$:
@@ 3.5 Training and Inference
4 Experiments
In this section, we conduct extensive experiments on four datasets to evaluate the robustness and generalization of the proposed Test-time Adaptive Hierarchical Co-enhanced Denoising Network (TAHCD) under multimodal noise, comparing it with eight representative reliable multimodal learning methods for low-quality data.
4.1 Experimental Settings
4.1.1 Datasets
Experiments are conducted on four commonly used multimodal datasets in previous works [13, 37, 39, 41, 5], including two multi-omics datasets. (1) BRCA: BRCA [22] is a breast invasive carcinoma dataset for PAM50 subtype classification, comprising 875 samples with features from three modalities: mRNA expression (mRNA), DNA methylation (meth), and miRNA expression (miRNA). Samples are categorized into five subtypes: Normal-like, Basal-like, HER2-enriched, Luminal A, and Luminal B. BRCA is available from The Cancer Genome Atlas (TCGA)  [^1] https://www.cancer.gov/aboutnci/organization/ccg/research/structuralgenomics/tcga. (2) ROSMAP: ROSMAP [24, 1, 10] comprises 351 samples from Alzheimer’s patients and normal controls, each with three modalities: mRNA expression (mRNA), DNA methylation (meth), and miRNA expression (miRNA). (3) CUB: The Caltech-UCSD Birds dataset [28] contains 11,788 samples across 200 bird categories, each with two modalities: bird images and corresponding textual descriptions. (4) UPMC FOOD101: The UPMC FOOD101 dataset [30] contains 90,704 samples from 101 food categories, each with an image obtained via Google Image search and a corresponding text description collected from uncontrolled environments, thus inherently containing noise.
4.1.2 Compared Methods
To validate the reliability of TAHCD under multimodal noise, we compare it with several representative reliable multimodal learning methods in our experiments, including multimodal dynamics (MD) [13], multi-level confidence learning (MLCLNet) [39], quality-aware multimodal fusion (QMF) [37], predictive dynamic fusion (PDF) [4], Semantic Invariance Learning (SMILE) [34], Seeking Proxy Point via Stable Feature Space (SPS) [33], Asymmetric Similarity Learning (ASL) [32], and Disentangled Noisy Correspondence Learning (DisNCL) [9]. The first four methods focus on learning high-confidence modality representations or fusing different modalities based on their qualities, and are therefore regarded as representative approaches for modality-specific noise removal. The remaining four methods emphasize learning cross-modality consistency and correspondence, and are thus treated as representative approaches for cross-modality noise removal.
4.1.3 Evaluation Metrics
BRCA & CUB & UMPC FOOD101
ROSMAP
4.1.4 Implementation Details
Training Details
Model Implementation Details
In TAHCD, the modality encoder $\phi_{x}^{m}$ is a fully connected network with input dimension equal to the modality dimension of the dataset and output dimension 256. The encoders $\phi_{\lambda}^{m}$, $\phi_{s}^{m}$, and $\phi_{c}^{m}$ are fully connected networks with input and output dimensions of 256 and a hidden layer of 512. The decoder $\psi^{m}$ is a fully connected network with input dimension 256 and output dimension equal to the modality dimension of the dataset.
Experimental Details
Two types of noise are involved in the experiments: modality-specific and cross-modality noise. Modality-specific noise is implemented following [12, 41, 4], with severity level $\epsilon$. Cross-modality noise follows [34, 9], where unaligned samples are introduced by shuffling the features of each modality for a random subset of samples with a proportion of $\eta$.
4.2 Comparison with State-of-the-Art Methods
To evaluate the robustness and generalization of TAHCD under multimodal noise, we conduct experiments on datasets corrupted by modality-specific and cross-modality noise [13, 4, 34, 9], with $\epsilon$ and $\eta$ denoting their respective severity levels.
4.2.1 Robustness
To evaluate the improved robustness of TAHCD against multimodal noise, we compare the classification performance of different methods on train and test sets that are both corrupted by the same noise, as reported in Table 1. MD, MLCLNet, QMF, and PDF learn reliable modality representations by estimating modality confidence but lack awareness of cross-modality alignment. Consequently, their performance is relatively stable under settings with only modality-specific noise ($\eta=0,\epsilon=5$), while their performance degrades significantly once cross-modality noise is introduced (all settings with $\eta=10\%$). SMILE, SPS, ASL, and DisNCL primarily focus on cross-modality alignment while paying limited attention to confidence of each modality. As a result, they exhibit relatively stable performance when only cross-modality noise is present ($\eta=10\%,\epsilon=0$), but their performance degrades noticeably once modality-specific noise is introduced (all settings with $\epsilon=5$). In contrast, TAHCD consistently achieves more stable performance than the compared methods under various noise conditions on all datasets, demonstrating superior robustness. This advantage primarily stems from its explicit modeling of both modality-specific and cross-modality noise and their joint removal across multiple levels, enabling more thorough noise suppression.